OpenStudy (anonymous):

I need help on solving the quadratic equations: x2 − 25 = 0   and   x2 = 2x + 15

OpenStudy (anonymous):

\[x^{2}-25 = (x-5)(x+5)\] so the solutions are -5 & +5. \[x^{2}-2x-15=(x-5)(x+3)\] so the solution is 5 & -3